OpenStudy (anonymous):

Choose the correct product of (5x − 3)(5x + 3). Is it 25x^2-9 ?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

Yep! Glad to see you're getting this down now :-D

OpenStudy (anonymous):

The middle two terms cancel out as you may have noticed (+15x - 15x = 0)

OpenStudy (asnaseer):

it is a specific instance of this factorisation:\[a^2-b^2=(a+b)(a-b)\]